We had a look at tonight injections and compared them with those of the previous night
The examined periods are:
17000 s starting from Dec 11, 01:00
17000 s starting from Dec 11, 22:20
We used a spectra resolution of 2000s.
The first dataset is polluted by a 0.1 Hz spaced comb, which is present in the Voltage and current monitors of the large coil amplifier as well as in the witness magnetometer (placed close to the WE tower). We suspect this might be related to the injected lines being multiples of 0.1 Hz. The problem with this is that Hrec itself is polluted by a 0.1 Hz comb, although it is not related with our injected noise. Infact, as shown in Figure 16 (example around the injected line 10.5Hz) the same 0.1 Hz comb was present also when not injecting (purple).
Therefore, in the second set of injections we moved the lins frequencies at **.*77 Hz. This time we noticed a somehow larger spectral noise "floor" in the Voltage, Current monitors and magnetometers as well, which resembles a comb spaced by 0.77 Hz. However, this turns out to be better, because no such comb is present in Hrec a priori.
Going through the injected lines one by one (Figures 1 to 15), it looks that we have achieved a good measurement of CF at most of them. Missing ones for now are 3.6, 4.6, 5.7 Hz, which might be recovered improving the analysis.
We do not have a good explanation for the extra noise produced by the amplifier, it seems related to a non linearity in the amplifier itself, since the spectrum of the injected noise (ENV_NOISE_MAG_WEB) does not show this features.